helpmeineedhelp:

Which equation can pair with 3x + 4y = 8 to create a consistent and independent system? 6x + 8y = 16 –3x – 4y = –6 6x – 3y = 2 –3x – 4y = –8

helpmeineedhelp:

Thank your for helping me

HaydenwH:

do u have to use substitutiin or sum else???

helpmeineedhelp:

Algebra is not my strong suit....

helpmeineedhelp:

Im not sure

HaydenwH:

ok one sec

helpmeineedhelp:

The topic is Solving Linear Systems Graphically

HaydenwH:

do you have any idea if u can eleminate one??

helpmeineedhelp:

Nope

helpmeineedhelp:

I have no idea how this works *sighs sadly*

helpmeineedhelp:

Lukaskosir it is the first one

HaydenwH:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @HaydenwH well u could prob start wth eliminating a,c \(\color{#0cbb34}{\text{End of Quote}}\) oop i was supposed to say b,d lmao

helpmeineedhelp:

lol

HaydenwH:

*feels dumb rn*

helpmeineedhelp:

whats your reasoning tho- why can we eliminate b and d?

Lukakosir:

Wait, was i right?

helpmeineedhelp:

Oh no

helpmeineedhelp:

I was saying that it is the first question

HaydenwH:

bc it doesnt fit into the equation

helpmeineedhelp:

Thank you hayden

Lukakosir:

Wait what happend

Lukakosir:

Why was i being warrned

HaydenwH:

your welcome, and nice try lukakosir

Lukakosir:

Huh?

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lukakosir Why was i being warrned \(\color{#0cbb34}{\text{End of Quote}}\) u gave a direct answer

HaydenwH:

bc u gave a direct answer

Lukakosir:

Bro sorry i dont know how to use this app

Jasonm:

try multiplying the problem you already have by a number to get an equivalent answer. like 3+3=6. multiply both 3s by 2, and they become 6, which changes your answer to 12.

Mercury:

Recall that a consistent *and* independent system has two equations with one solution only. That means they only intersect once. Therefore, the two equations must have different y-intercepts *and* slopes. To help you out: if you multiply 3x + 4y = 8 by 2, you get 6x + 8y = 16 if you multiply 3x + 4y = 8 by (-1) you get -3x - 4y = -8 neither of these answers works. if you multiply an equation by a constant, you get the same equation. that's not what were looking for. between the other two choices: which one is an equation with a different slope *and* y-intercept from the original?

helpmeineedhelp:

Of which two answers?

helpmeineedhelp:

Im trying to find the slope and intercept and see if it's different from that of the original equation, so which two equations did you eliminate?

Mercury:

this is the original equation: 3x + 4y = 8 I eliminated 6x + 8y = 16 and -3x - 4y = -8 because they're just multiples of the original equation between –3x – 4y = –6 and 6x – 3y = 2, see which one has a different slope and y-intercept compared to 3x + 4y = 8 if you're having trouble: solve all three equations for y

helpmeineedhelp:

Thank you

helpmeineedhelp:

So it's C, 6x-3y=2, right? Because it has a different slope and y intercept than 3x+4y=8?

Mercury:

yup good

helpmeineedhelp:

Thank you for your help